ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 32 Kohr, Gabriela, Babeş-Bolyai University, Cluj-Napoca, Romania Univalent mappings of several complex variables In this work we will present some personal contributions in the domain of the geometric function theory of several complex variables. More precisely, we will discuss several problems about some subclasses of biholomomorphic mappings from pseudoconvex domains in C^{n} into C^{n}. Of particular interest is the case of starlike, convex and spirallike mappings on the unit ball and other domains in C^{n}. For these classes we point out some interesting problems and examples. Reference: G. Kohr, P. Liczberski, Univalent Mappings of Several Complex Variables, Cluj University Press, 1998.